Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding

Fabrication

Total

Estimated total machine-hours used

2,500

1,500

4,000

Estimated total fixed manufacturing overhead

$

11,250

$

15,750

$

27,000

Estimated variable manufacturing overhead per machine-hour

$

1.90

$

2.70

Job P

Job Q

Direct materials

$

18,000

$

10,500

Direct labor cost

$

25,000

$

9,500

Actual machine-hours used:

Molding

2,200

1,300

Fabrication

1,100

1,400

Total

3,300

2,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month. assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

9. What were the company’s predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your answers to 2 decimal places.)

10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.)
Answer #1
9
Predetermined overhead rates :
Molding Department 6.40 =1.90+(11250/2500)
Fabrication Department 13.20 =2.70+(15750/1500)
10
Manufacturing overhead applied:
Job P 14080 =2200*6.40
Job Q 8320 =1300*6.40
11
Manufacturing overhead applied:
Job P 14520 =1100*13.20
Job Q 18480 =1400*13.20
12
Job P
Direct materials 18000
Direct labor cost 25000
Manufacturing overhead applied 28600 =14080+14520
Total manufacturing cost Job P 71600
Divide by units 20
Unit product cost Job P 3580